The number of tickets sold each day for an upcoming performance of Handel's Messiah is given by N(x)=-0.5x^2+14x+10, where x is the number of days since the concert was first announced. When will daily ticket sales peak and how many tickets will be sold that day?
Answer by scott8148(6628)

wow..the "Hallelujah" chorus...one of my favorites

the peak (maximum) is on the axis of symmerty
__ the general equation is x=-b/2a
__ in this case, x=-14/[2(-0.5)] __ x=-14/(-1) __ x=14
__ so, sales peak on the 14th day

to find the number of tickets, substitute into the equation
__ N=-.05(14)^2+14(14)+10 __ N=108